### v1

Abstract
Explainable Artificial Intelligence (XAI) techniques, such as SHapley Additive exPlanations (SHAP), have become essential tools for interpreting complex ensemble tree-based models, especially in high-stakes domains such as healthcare analytics. However, SHAP values are usually treated as point estimates, which disregards the inherent and ubiquitous uncertainty in predictive models and data. This uncertainty has two primary sources: aleatoric and epistemic. The aleatoric uncertainty, which reflects the irreducible noise in the data. The epistemic uncertainty, which arises from a lack of data. In this work, we propose an approach for decomposing uncertainty in SHAP values into aleatoric, epistemic, and entanglement components. This approach integrates Dempster-Shafer evidence theory and hypothesis sampling via Dirichlet processes over tree ensembles. We validate the method across three real-world use cases with descriptive statistical analyses that provide insight into the nature of epistemic uncertainty embedded in SHAP explanations. The experimentations enable to provide more comprehensive understanding of the reliability and interpretability of SHAP-based attributions. This understanding can guide the development of robust decision-making processes and the refinement of models in high-stakes applications. Through our experiments with multiple datasets, we concluded that features with the highest SHAP values are not necessarily the most stable. This epistemic uncertainty can be reduced through better, more representative data and following appropriate or case-desired model development techniques. Tree-based models, especially bagging, facilitate the effective quantification of epistemic uncertainty.
keywords:
[…]
1 Introduction
Machine learning (ML) [1] is a key part of improving healthcare analytics such as resource planing, disease diagnosis, prognosis, and risk stratification [2, 3, 4]. However powerful, uncertainty is inherent and ubiquitous in machine learning (ML) models because their predictions are affected by noisy data, model limitations, and unseen scenarios. To address this challenge, some of the most widely used tools are ensemble tree-based models [5], which help in managing and quantifying uncertainty in predictions. They are highly accurate, interpretable, and efficient with structured data, resulting in lower computational demand. Unlike deep neural networks, which require large amounts of unstructured data such as images and text, they have lower computational requirements and are more interpretable [6, 7]. These include Random Forest (RF) [8], Gradient Boosting Machines (GBM), and Extreme Gradient Boosting (XGBoost) [9]. These models are robust against noise, and can handle large, complicated data sets, which are common in healthcare [10]. Ensemble tree approaches are different from traditional machine learning models [11] as they can efficiently capture complex, nonlinear relationships and slight interactions among the features [12]. This leads to highly accurate and generalizable predictions. Ensemble models have many advantages. They combine the strengths of multiple base learners which reduces overfitting, improves stability, and enhances the model’s ability to generalize to unseen data unlike traditional ML models [7]. It has been shown that using a group of classifiers to make predictions outperforms using individual classifiers to predict heart disease [10, 13]. This is important for making reliable decisions, which makes them very useful in healthcare.Despite their strengths, ensemble tree-based models have two main problems. First, they are difficult to interpret, particularly when there are large numbers of constituent trees and features [14, 15, 16]. This “black box” nature, avoiding them to be the first choice in the model selection. To address this, explainable AI (XAI) techniques such as SHapley Additive exPlanations (SHAP) or Shapley values [17], which are rooted in cooperative game theory, have emerged as a principled framework for attributing the contribution of each feature to individual predictions in ML models. However, calculating Shapley values can be difficult for complex models. Fortunately, a couple of new, efficient methods for calculating them for certain types of models have recently emerged. TreeSHAP [18, 19], a fast and exact method for calculating Shapley values in tree-based models like decision trees, RFs, XGBoost, have been introduced. TreeSHAP uses the natural structure of decision trees to make predictions that are much faster and easier to understand than those of other methods. This is helpful for explaining results with large groups of data, which is important in fields where understanding predictions is paramount, for instance in healthcare or finance. SHAP values provide a clear framework for determining how each feature contributes to individual predictions. They offer insight into the decision-making process behind complex ensemble models. SHAP and similar methods not only encourage trust, but also make it easier to use advanced machine learning (ML) models in healthcare by making them easier to comprehend.
Recent research has identified several factors that improve SHAP values. These factors include misattribution of feature importance, reliance on the assumption of feature independence, lack of causal or contextual understanding, computational inefficiency, and risk of misinterpretation. To address these issues, alternative attribution methods and error quantification techniques, such as Normalized Movement Rate (NMR) and Modified Index Position (MIP), have been proposed to handle feature collinearity [20]. New SHAP variants have also been developed, with the aim of improving efficiency and interpretability. Latent SHAP [21, 19] extends SHAP by enabling explanations in human-interpretable domains without requiring invertible transformations, making it suitable for high-dimensional or non-invertible data and capturing correlations among features. Kernel SHAP [22], the most versatile black-box SHAP explainer, uses weighted linear regression to approximate Shapley values and is valued for its generality. However, it is slower and assumes feature independence, which can limit accuracy in correlated data. Muschalik et al. [23] introduce methods for efficiently computing higher-order Shapley interactions in tree ensembles. This enables richer, more granular explanations of feature interactions than standard SHAP, with significant computational advantages for large or complex models. These advancements collectively enhance the reliability, interpretability, and practicality of SHAP-based explanations in ML. Other advancements include integrating causal and contextual information, as well as creating more computationally efficient SHAP variants, such as CF-SHAP and FF-SHAP [24].
However, SHAP is a point-estimate method, which contributes to the uncertainty of the explanations it produces. This opens up new dimensions for studying and quantifying uncertainties in XAI, which is an important step forward in the field [25, 26, 27]. Current implementations, as discussed before, treat SHAP or TreeSHAP values with respect to tree-based ML models as point estimates. However, this approach ignores the individual contributions of epistemic uncertainty arising from variability in model training. It focuses on overall uncertainty, comprising aleatoric and epistemic uncertainty. Aleatoric uncertainty refers to the inherent randomness or noise in the data that cannot be reduced by collecting more information. Epistemic uncertainty, on the other hand, arises from a lack of knowledge or data about the model or process. This omission poses critical risks in high-stakes domains. For instance, medical diagnostics using XGBoost may yield identical SHAP values across hospitals despite shifts in regional data distribution. Similarly, financial risk models may exhibit unstable feature attributions during market volatility [28, 29]. Random forest (SHAP) models used for predictive health monitoring [30, 31] may appear similar or stable when applied to hospitals with different demographics or disease prevalence rates but SHAP values often exhibit bias toward features with higher cardinality or entropy. This bias can overstate or understate the importance of these features when patient populations change. Quantifying epistemic uncertainty in SHAP values is necessary because people trust model explanations for decision support, especially in high-stakes domains such as healthcare and finance [32, 33, 25]. Contemporary methods use techniques such as bootstrap sampling to estimate the uncertainty of SHAP attributions and generate confidence intervals for the importance of each feature. Now, variants of SHAP enable users to evaluate the reliability of feature contributions instead of relying exclusively on point estimates. These methods allow practitioners to more accurately evaluate the stability of explanations, identify features or contexts in which explanations are less reliable, and improve model transparency in situations involving shifting or uncertain data.
The prevailing methods of uncertainty quantification (UQ) predominantly focus on assessing predictive uncertainty by integrating aleatoric and epistemic uncertainty. In this research, we introduce a framework that:
  1. 1.
    Decomposes SHAP variance into aleatoric and epistemic components
  2. 2.
    Leverages belief functions and Dirichlet processes for hypothesis space sampling
  3. 3.
    Provides computationally tractable epistemic uncertainty intervals for feature attributions.
2 Background
2.1 SHAP
The Shapley values are predicated on the strong foundation of cooperative game theory. For a set of players $N$ and a value function $v$, the Shapley value $\phi_{i}$ for player (or feature) $i$ is defined as:
$\phi_{i}(N,v)=\frac{1}{|N|!}\sum_{\text{all orderings }R}\left[v(P_{i}^{R}\cup\{i\})-v(P_{i}^{R})\right]$
where:
  • $R$ is a permutation (ordering) of all players.
  • $P^{R}_{i}$ is the set of players that precede $i$ in ordering $R$.
  • $v(S)$ is the value (e.g., model output) associated with subset $S\subseteq N$ of players.
  • Marginal Contribution: $v(S\cup\{i\})-v(S)$
  • Averaging: Weighted by the number of ways each coalition can be formed in all possible player orderings.
  • Efficiency: $\sum_{i\in N}\phi_{i}=v(N)$, ensuring that the total value is fairly distributed among all players.
Shapley values enable the quantification and interpretation of feature contributions in ML. Research shows SHAP as one of the most interpretable methods in ML, providing insight into complex healthcare ML models. It is a model-agnostic interpretability tool used extensively in healthcare. [34, 35, 36, 37]. Furthermore, it has been used in fields such as predicting breast cancer risk, elucidating model predictions, diagnosing biomarkers, and analyzing survival, particularly with tree-based machine learning models [38, 39, 10]. SHAP has been used to interpret machine learning models that predict cancer risk. For example, it has identified age and family history as key predictors of breast cancer risk [38]. Furthermore, SHAP has improved the interpretability of models for other chronic diseases. It has been used to examine machine learning models for smoking and drinking habits, using lifestyle data, blood test results, and wearable sensor readings. This has facilitated the interpretation of key influencing features and enhanced transparency for potential use in personalized healthcare [40]. In use cases involving imaging or clinical data, SHAP reveals the importance of diagnostic features. For example, SHAP emphasizes the texture and morphology of tumors in breast cancer mammography and the key variables in detecting chronic diseases[41]. The field of radiology stands to benefit from SHAP because it facilitates the use of AI models to interpret imaging scans for abnormalities, such as lung nodules or diabetic retinopathy [42]. This development has the potential to improve both diagnostic accuracy and patient trust. In the context of retinoblastoma diagnosis, SHAP was employed to generate local and global interpretations, highlighting specific regions and features in fundus images that substantially influence the model’s predictions  [43]. SHAP was also used in deep learning models to identify image features that facilitate early cancer detection and enhance the interpretability of automated histopathology analyses [44]. Applying SHAP to deep learning (DL) models in medical image analysis provides clinicians with visual interpretations of model predictions. This improves the understanding and validation of automated diagnoses in various imaging tasks [45]. Recent frameworks continue to adopt SHAP as a technique for improving explainability in healthcare [46].
2.2 Uncertainty in XAI
Background and Approaches to Uncertainty Quantification in XAI:
There is a growing interest and there have been recent advances which focus on developing methods for uncertainty quantification in XAI [47, 48]. These methods focus on communicating the uncertainty associated with the interpretations, which is necessary for the wider adoption of AI in high-stakes scenarios. The quantification of the uncertainty related to the interpretations involves the study of the change in interpretation when the input data or the model parameters are changed. One of the recent works introduces a framework that models the interpretations as a function $e_{\theta}(x,f)$ where for a model $f$, an instance $x$, and explanation parameters $\theta$, the explanation $e_{\theta}(x,f)$ quantifies each feature’s contribution to the prediction  [49]. The function allows researchers to follow the uncertainty from the inputs and the model with the help of interpretations. Methods like this one frequently use empirical and analytical estimations. The former often include Monte Carlo simulations techniques that enable researchers to obtain multiple different version of the input or the model, allowing them to study the variance of the model output and corresponding interpretations. The latter, focus on how small changes in the inputs or the model parameters affect the interpretations, allowing the creation of a co-variance matrix that quantifies the uncertainty. Another recent work [49] introduces the Mean Uncertainty in the Explanations (MUE) metric which summarizes the overall uncertainty by normalizing the trace of the interpretations’ co-variance matrix. The method also enables researchers to directly compare the uncertainty values across different methods and models. However, most of the studies mentioned earlier show that XAI methods are only point estimates. This means that they can identify important features, but they don’t show how reliable the interpretability results are when different inputs and model parameters are used [50].
SHAP Specific Uncertainty Quantification:
Some of the works have made significant achievements in the terms of quantifying the uncertainty of SHAP values by calculating the confidence intervals and distributions. This is crucial in decision-making in healthcare [51]. The way traditional SHAP scores are calculated is limited because they rely on input data that is either well-specified or estimated accurately [52]. The standard SHAP framework requires knowledge of the underlying probability distribution, which is often either unknown or estimated from a small number of samples. This can lead to unstable or misleading feature importance estimates. A recent framework [26], calculates the SHAP score as a function over a range of possible distributions. The approach provides tight intervals for feature importance and shows that feature rankings can be very sensitive to distributional assumptions. The framework also shows that related decision problems can’t be solved using computers. In other words, these problems are NP-complete. These problems include determining if a feature SHAP score can be higher than a threshold or always outperform another feature, even for decision trees. Studies have shown that SHAP intervals can be quite wide when there’s uncertainty about the data distribution [26]. However, as more data becomes available, these intervals become more stable.
Information Theory Approaches and Other Alternatives:
Researchers have come up with new ways to understand the predictions and the uncertainty of models. These new ways use information theory to explain the predictions and uncertainty. They also help reduce uncertainty in certain features and provide efficient algorithms and methods for making inferences in the real world. These research areas show how important it is to understand the assumptions about data distribution. This is important for making sure that the features used in explainable machine learning can be trusted and understood. In healthcare, incorporating uncertainty quantification has allowed researchers to develop models more reliable and transparent [53]. It is very important to combine uncertainty quantification with XAI in high-stakes application, especially when the people who know the most about the domain utilize both the model predictions and explanations to make decisions. Recent research [25, 54, 32] on quantifying uncertainty in SHAP has improved the interpretability of ML models. The research by Cohen et al. [32] and Watson et al. [25] are significant in terms of extending XAI methods based on SHAP values to better quantify and explain uncertainty in model predictions. Cohen et al. introduce efficient methods and visual tools for measuring uncertainty in stochastic SHAP explanations, making them more applicable to real-time or high-stakes scenarios. Watson et al. use information theory to expand SHAP values and quantify the predictive uncertainty, offering formal reliability guarantees and scalable algorithms for practical use. But, both the approaches are sensitive to data and model quality. They primarily address uncertainty from estimator sampling. The challenges in terms of assessing broader sources of uncertainty. Overall, these significant developments help make AI explanations more transparent and trustworthy, especially with regard to uncertainty. However, further research is needed to study uncertainty arising from models and data.
Although these works enable users to evaluate feature importance and model attribution confidence, additional dimensions must be addressed. Current methods include uncertainty aggregation and assign non-zero importance to irrelevant features. They are also sensitive to data and model biases, computationally intensive, and may produce unreliable attributions under model instability. Additionally, they lead to misleading interpretations in high-stakes scenarios and cannot distinguish between aleatoric and epistemic uncertainty. These methods are also vulnerable to adversarial manipulation. In the context of healthcare, SHAP values may lead domain professionals to overconfidently rely on feature importance when making treatment decisions based on model explanations. They may ignore epistemic uncertainty because data can be limited, biased, or heterogeneous. Furthermore, SHAP does not represent aleatoric uncertainty due to noisy or ambiguous input data. Misleading explanations can also occur if the model is biased, the data is collinear, or the underlying relationships are not well captured. Decomposing SHAP uncertainty into epistemic and aleatoric components can address these limitations [55]. This allows for targeted model improvement and enhanced, actionable explanations. However, as commonly implemented, SHAP values do not account for epistemic uncertainty, which arises from variability in model training. This limitation arises from SHAP’s design as a post hoc explanation tool for individual predictions rather than as a method for quantifying uncertainty.
3 Problem Formulation
We can introduce a theorem as follows:
Theorem 1.
For any tree ensemble model $f$, the point estimate SHAP value $\phi_{i}$ lacks a measure of variance V($\phi_{i}$|f, D) over possible training datasets $D\sim P_{\text{data}}$. This violates the reliability axiom for explainability in high-risk AI systems [56].
This serves as the the motivation to decompose the SHAP variance into aleatoric, epistemic, and covariance terms as follows:
3.1 SHAP Variance Decomposition
The conventional uncertainty quantification framework posits that aleatoric uncertainty stems from inherent data noise and that epistemic uncertainty stems from model ignorance. However, empirical evidence challenges this distinction by demonstrating that, under shifts in the data distribution or model misspecification, aleatoric and epistemic uncertainties become intertwined. Bootstrap ensembles and deep ensemble methods show that, as epistemic uncertainty increases, estimates of aleatoric uncertainty can decrease, leading to systematic bias in model predictions [57, 58, 59]. For any feature $i$ and instance $\mathbf{x}$, the total variance of SHAP values $\phi_{i}(\mathbf{x})$ over possible training datasets $D\sim P_{\text{data}}$ and tree ensemble models $f$ decomposes as:
$\underbrace{\text{Var}_{D,f}(\phi_{i})}_{\text{Total}}=\underbrace{\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|f)]}_{\text{Aleatoric}}+\underbrace{\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])}_{\text{Epistemic}}+\underbrace{2\cdot\mathcal{C}(f,D)}_{\text{Entanglement}}$ (1)
where $\mathcal{C}(f,D)=\text{Cov}_{P(f,D)}\left(\mathbb{E}_{D}[\phi_{i}|f],\text{Var}_{f}(\phi_{i}|D)\right)$, and $\phi_{i}(x)$ The terms in the equations can be defined as follows:
  1. 1.
    Aleatoric Uncertainty ($\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]$):
    • Variance from model stochasticity (tree structure randomization) for fixed $D$
    • For tree ensembles: reflects variability due to bootstrap sampling and feature randomization
  2. 2.
    Epistemic Uncertainty ($\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$):
    • Variance from data sampling (different $D$ yield different mean SHAP values)
    • Measures sensitivity to training data composition
  3. 3.
    Entanglement Term ($\mathcal{C}(f,D)$):
    • Covariance between mean SHAP ($\mathbb{E}[\phi_{i}|f]$) and SHAP variability ($\text{Var}(\phi_{i}|D)$)
    • Non-zero when: Models producing higher mean $|\phi_{i}|$ exhibit higher variance (common in tree ensembles due to node splitting)
    • The covariance indicates whether features with higher average absolute SHAP values also tend to have greater variance in their SHAP values. This covariance is particularly non-zero in tree ensemble models because of the nature of node splitting.
Proof
Let:
  • $\phi_{i}(\mathbf{x}|f,D)$: SHAP value for feature $i$ on instance $\mathbf{x}$ given model $f$ trained on dataset $D$
  • $f\sim P(f|D)$: Tree ensemble model distribution (via bootstrap/randomization in training)
  • $D\sim P_{\text{data}}$: Data distribution
  • $P(f,D)=P(f|D)P(D)$: Joint distribution
We have assumed the following assumptions:
  1. 1.
    Model-Dataset Separability: $P(f,D)=P(f|D)P(D)$ (standard ML training)
  2. 2.
    Finite Variance: $\text{Var}(\phi_{i}|f)$ and $\text{Var}(\mathbb{E}[\phi_{i}|D])$ exist $\forall f,D$
  3. 3.
    SHAP Linearity: $\phi_{i}$ is linear in tree outputs (holds for TreeSHAP)
Law of Total Variance
Apply the law of total variance (first decomposition) [60] conditioned on $f$:
$\text{Var}_{D,f}(\phi_{i})=\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$ (2)
This gives the standard aleatoric (first term) and epistemic (second term) decomposition, but ignores the model-data dependency.
Refinement for Entanglement
The term $\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]$ is decomposed by conditioning on $D$:
$\begin{split}\mathbb{E}_{f}[\mathrm{Var}_{D}(\phi_{i}|f)]&=\mathbb{E}_{D}[\mathrm{Var}_{f}(\phi_{i}|D)]\\ &\quad+\left(\mathbb{E}_{f}[\mathrm{Var}_{D}(\phi_{i}|f)]-\mathbb{E}_{D}[\mathrm{Var}_{f}(\phi_{i}|D)]\right)\end{split}$ (3)
The excess term arises from non-commutativity of expectations due to $P(f,D)\neq P(f)P(D)$.
Covariance Identification
The entanglement term emerges from:
$\mathcal{C}(f,D)=\text{Cov}\left(\mathbb{E}_{D}[\phi_{i}|f],\text{Var}_{f}(\phi_{i}|D)\right)$ (4)
Derivation:
  1. 1.
    Expand $\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]$ using iterated expectation:
    $\begin{split}\mathbb{E}_{f}[\text{Var}_{D}(\phi_{i}|f)]=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\\ \text{Cov}\left(\mathbb{E}_{D}[\phi_{i}|f],\text{Var}_{f}(\phi_{i}|D)\right)\end{split}$ (5)
  2. 2.
    Substitute into total variance:
    $\text{Var}_{D,f}(\phi_{i})=\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]+\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])+2\mathcal{C}(f,D)$ (6)
Special Case: Tree Ensembles
For Random Forests with $B$ trees trained on bootstrap samples $\{D_{b}\}$:
  1. 1.
    Aleatoric Term:
    $\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]\approx\frac{1}{B}\sum_{b=1}^{B}\text{Var}_{T\in\mathcal{T}_{b}}(\phi_{i}^{(T)})$ (7)
    where $\mathcal{T}_{b}$ are trees trained on $D_{b}$.
  2. 2.
    Epistemic Term:
    $\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])\approx\text{Var}_{b}\left(\frac{1}{|\mathcal{T}_{b}|}\sum_{T\in\mathcal{T}_{b}}\phi_{i}^{(T)}\right)$ (8)
  3. 3.
    Entanglement Term:
    $\mathcal{C}(f,D)\propto\sum_{b=1}^{B}\left(\bar{\phi}_{i}^{(b)}-\bar{\phi}_{i}\right)\left(\sigma_{i}^{(b)2}-\bar{\sigma}_{i}^{2}\right)$ (9)
    where:
    • $\bar{\phi}_{i}^{(b)}=$ mean SHAP for trees in bootstrap $b$
    • $\sigma_{i}^{(b)2}=$ SHAP variance for trees in bootstrap $b$
The UbiQTree estimator approximates this decomposition via:
  1. 1.
    Dirichlet Sampling: Simulates $P(f|D)$ by weighting trees via OOB performance
  2. 2.
    Variance Components:
    • Aleatoric: Variance of SHAP across trees within each weighted sample
    • Epistemic: Variance of mean SHAP across samples
    • Entanglement: Covariance between sample means and variances
The proof enables us to list out the facts that:
  1. 1.
    SHAP variance decomposition requires accounting for model-data entanglement in tree ensembles
  2. 2.
    The entanglement term $\mathcal{C}(f,D)$ is non-negligible when:
    • Feature importance correlates with SHAP variability (common in high-gain features)
    • Data distributions induce model instability (e.g., rare categories)
  3. 3.
    E-SHAP’s Dirichlet-weighted sampling preserves this covariance structure, unlike bootstrap methods that assume $P(f,D)\approx P(f)P(D)$
The decomposition enables precise uncertainty attribution in feature importance analysis, critical for high-stakes applications.
3.2 Evidence Theory: Tree Ensembles
The Dempster-Shafer theory (DST) is a mathematical framework for reasoning under uncertainty, particularly when evidence is incomplete, imprecise, or conflicting. DST assigns belief masses to sets or intervals of possible outcomes, allowing for the explicit representation of ignorance and epistemic uncertainty. Key DST concepts include belief mass ($m$), belief ($Bel$), plausibility ($Pl$), and ignorance. DST is widely used in artificial intelligence, sensor fusion, medical diagnostics, risk assessment, and autonomous systems, where managing uncertainty and combining evidence from multiple sources is critical. It provides a systematic and flexible approach to uncertainty, enabling AI and decision systems to model ignorance and combine evidence in ways that classical probability theory cannot. DST is a valuable tool for managing uncertainty and combining evidence in AI and decision systems. [61, 53].
Dempster-Shafer Representation
For a tree ensemble with $K$ trees, the Basic Probability Assignment (BPA) for SHAP value $\phi_{i}$ belonging to interval $A\subseteq\mathbb{R}$ is:
$m(A)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{I}\left(\phi_{i}^{(k)}\in A\right)$ (10)
where $\phi_{i}^{(k)}$ is the SHAP value from tree $T_{k}$. The Belief and Plausibility functions satisfy:
$\text{Bel}(A)=\sum_{B\subseteq A}m(B),\quad\text{Pl}(A)=\sum_{B\cap A\neq\emptyset}m(B)$ (11)
Proof: BPA Construction Each tree represents an independent evidence source. The BPA is the proportion of trees supporting interval $A$, satisfying:
  • $m(\emptyset)=0$ (impossible event)
  • $\sum_{A\subseteq\mathbb{R}}m(A)=1$ (normalization)
Belief Function:
For nested intervals $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{n}$:
$\text{Bel}(A_{n})=\sum_{j=1}^{n}m(A_{j})\quad\text{(consonant structure)}$ (12)
This follows from the definition of Belief as the total evidence supporting $A$.
Plausibility Bound:
For conflicting explanations (e.g., positive vs. negative impact):
$\text{Pl}(A)-\text{Bel}(A)=1-\sum_{B\subseteq A}m(B)-\sum_{B\subseteq A^{c}}m(B)$ (13)
where $A^{c}$ is the complement. This quantifies the probability mass assigned to sets overlapping both $A$ and $A^{c}$.
Tree Ensemble Specialization:
Since trees are exchangeable:
$\lim_{K\to\infty}\text{Bel}(A)=\mathbb{P}(\phi_{i}\in A)$ (14)
By the Law of Large Numbers, Belief converges to the true probability.
Conflict Measure: The explanation conflict for feature $i$ is:
$\mathcal{C}_{i}=\sup_{A\subseteq\mathbb{R}}\left[\text{Pl}(A)-\text{Bel}(A)\right]$ (15)
which measures the maximum ambiguity in SHAP assignments.
3.3 Uncertainty Theory: Application to SHAP
Uncertainty theory by Liu et al. [62, 63] is a mathematical framework designed to address epistemic uncertainty arising from incomplete knowledge, small sample sizes, or reliance on expert judgment. The theory is based on four axioms: normality, monotonicity, self-duality, and countable subadditivity. The central concept is the uncertainty distribution, denoted by the symbol $\Gamma\colon\mathbb{R}\to[0,1]$, which quantifies the degree of belief that a variable takes on values less than or equal to $\phi_{i}$. This distribution is characterized by a value of 0 for implausible values and a value of 1 for fully plausible values. It is also monotonically increasing as values become more plausible. Uncertainty distributions model subjective confidence rather than frequency or likelihood. This makes them useful in situations with limited or non-statistical data. When applied to SHAP or feature attribution in AI, an uncertainty distribution can represent confidence in a feature’s attribution magnitude. This allows practitioners to explicitly model and quantify their uncertainty about the importance of each feature, especially when data is scarce or unreliable. Entropy minimization can guide optimal data acquisition, thereby improving model interpretability and reliability [64, 62, 63, 65, 66].
Theorem 2 (Uncertainty Distribution).
The uncertainty distribution $\Gamma:\mathbb{R}\to[0,1]$ for SHAP value $\phi_{i}$ satisfies:
  1. 1.
    $\Gamma(c)=0$ for $c<\min_{k}\phi_{i}^{(k)}$
  2. 2.
    $\Gamma(c)=1$ for $c\geq\max_{k}\phi_{i}^{(k)}$
  3. 3.
    $\Gamma$ is monotonically increasing
Boundary Conditions:
By definition, implausible values (outside $\min\phi,\max\phi]$) have $\Gamma(c)=0$, and fully plausible values ($c\geq\max\phi$) have $\Gamma(c)=1$.
Monotonicity:
For any $c_{1}<c_{2}$:
$\{k:\phi_{i}^{(k)}\leq c_{1}\}\subseteq\{k:\phi_{i}^{(k)}\leq c_{2}\}$ (16)
Thus $\Gamma(c_{1})\leq\Gamma(c_{2})$ by set inclusion.
Entropy Minimization:
The uncertainty entropy is:
$H(\Gamma)=-\int_{-\infty}^{\infty}\gamma(c)\log\gamma(c)dc$ (17)
where $\gamma(c)=d\Gamma/dc$. Data acquisition minimizes $H(\Gamma)$ by:
$\mathbf{x}^{*}=\arg\min_{\mathbf{x}}\mathbb{E}_{y|\mathbf{x}}\left[H(\Gamma_{\mathcal{D}\cup(\mathbf{x},y)})\right]$ (18)
This follows from the information gain principle.
Lemma 1 (Optimal Acquisition).
When acquiring data for feature $j$, the uncertainty entropy decreases as:
$\Delta H\propto-\text{Cov}\left(\phi_{j},\frac{\partial\phi_{i}}{\partial\theta}\right)$ (19)
where $\theta$ is the model parameter space.
3.4 Dirichlet Process Hypothesis Sampling
Dirichlet processes (DPs) are key to Bayesian nonparametric modeling. They allow for flexible inference over distributions with unknown and potentially infinite underlying clusters. DPs are useful for modeling uncertainty in complex spaces like SHAP values and their clusters across tree ensembles. Each sample from a DP is a discrete probability distribution. DPs are parameterized by a concentration parameter, $\alpha$, and a base distribution, $G_{0}$. In mixture models, DPs allow for an unbounded number of mixture components, thereby adapting model complexity to the data. In tree ensembles, each tree is a hypothesis about feature attributions. By modeling the distribution of SHAP values across trees with a DP mixture model, one can cluster SHAP values without specifying the number of clusters or modes in advance. This method captures both diversity and epistemic uncertainty in feature attributions due to model variability and quantifies uncertainty by examining the posterior distribution of clusters or modes of SHAP values. This provides richer uncertainty estimates than standard bootstrap or ensemble variance methods. DPs have several advantages over parametric and bootstrap methods. First, they avoid the need to fix the number of clusters or modes in advance. DPs adapt to model complexity as more data or trees are considered. DPs mitigate under- or overfitting and provide a more nuanced, probabilistic view of uncertainty in SHAP attributions. DPs are widely used in machine learning for clustering and mixture models where the number of components is unknown. Such an approach allows for a richer and more flexible quantification of epistemic uncertainty in SHAP attributions by leveraging the full power of Bayesian nonparametrics [67, 68, 69, 70, 71, 72].
Theorem 3 (Constructing the Dirichlet Process).
The posterior over tree ensembles is given by:
$G\sim\text{DP}(\alpha,G_{0}),\quad G_{0}=\sum_{k=1}^{K}w_{k}\delta_{T_{k}},\quad w_{k}=\frac{\text{OOB-AUC}_{k}}{\sum_{j}\text{OOB-AUC}_{j}}$ (20)
Base Measure:
$G_{0}$ is a discrete measure weighted by out-of-bag (OOB) accuracy, satisfying $\int dG_{0}=1$.
Dirichlet Process:
For any partition $(B_{1},\dots,B_{m})$ of the tree space:
$(G(B_{1}),\dots,G(B_{m}))\sim\text{Dirichlet}(\alpha G_{0}(B_{1}),\dots,\alpha G_{0}(B_{m}))$ (21)
Concentration Parameter:
  • As $\alpha\to 0$: $G$ concentrates on $\max(w_{k})$ trees
  • As $\alpha\to\infty$: $G\to G_{0}$ (base measure)
SHAP Distribution:
The SHAP value distribution is:
$\mathbb{F}_{i}(A)=\int\phi_{i}(T)dG(T)$ (22)
With first moment:
$\mathbb{E}[\phi_{i}]=\sum_{k=1}^{K}\pi_{k}\phi_{i}^{(k)},\quad\pi\sim\text{Dirichlet}(\alpha\mathbf{w})$ (23)
Theorem 4 (Convergence).
As $K\to\infty$, the SHAP distribution converges:
$\mathbb{F}_{i}\xrightarrow{d}\mathcal{GP}\left(m(\mathbf{x}),\kappa(\mathbf{x},\mathbf{x}^{\prime})\right)$ (24)
where $m(\cdot)$ is the mean function and $\kappa(\cdot,\cdot)$ the covariance kernel.
  1. 1.
    By the de Finetti theorem, infinite exchangeable trees induce a Gaussian process [73, 74, 75, 76].
  2. 2.
    The Dirichlet process is the de Finetti measure for Pólya sequences [77].
  3. 3.
    SHAP values are continuous linear operators, preserving convergence [17].
4 Methodology
Our approach integrates three complementary theoretical frameworks to facilitate the decomposition and quantification of uncertainty in SHAP values: Dirichlet process hypothesis sampling, Liu’s uncertainty theory, and Dempster–Shafer theory. This integrated approach explicitly models the entanglement between aleatoric and epistemic uncertainties in feature attribution, overcoming the drawbacks of traditional uncertainty quantification. The framework allows for a thorough examination of sources of uncertainty(Algorithm: 5).
Evidence Theory for SHAP Uncertainty
Dempster-Shafer evidence theory provides a formal mechanism to represent ambiguity in SHAP distributions through belief functions. For a tree ensemble with $K$ trees, we construct a Basic Probability Assignment (BPA) over SHAP intervals $A\subseteq\mathbb{R}$ (Equation: 10). where $\phi_{i}^{(k)}$ denotes the SHAP value from tree $T_{k}$ (Algorithm: 2). The belief $\text{Bel}(A)$ and plausibility $\text{Pl}(A)$ functions then quantify the minimum and maximum support for interval $A$, respectively. The physical interpretation reveals that $\text{Bel}(A)$ represents conservative certainty (e.g., "SHAP lies in [1,1][-1,1] with $\geq$80% confidence"), while the conflict measure $\mathcal{C}_{i}=\sup_{A}[\text{Pl}(A)-\text{Bel}(A)]$ captures explanation ambiguity. High conflict triggers human verification in critical applications, and the BPA dispersion directly measures aleatoric uncertainty. This approach links to the SHAP variance decomposition by mapping belief/plausibility bounds to epistemic uncertainty ($\text{Var}_{f}(\mathbb{E}_{D}[\phi_{i}|f])$) and the conflict term to entanglement ($\mathcal{C}(f,D)$).
Uncertainty Theory for SHAP Uncertainty
Liu et al. uncertainty theory models epistemic uncertainty through the uncertainty distribution $\Gamma(c)=\mathbb{P}(\phi_{i}\leq c)$, bounded by $\min_{k}\phi_{i}^{(k)},\max_{k}\phi_{i}^{(k)}]$. The distribution’s shape provides intuitive study of the uncertainty : a steep $\Gamma$ indicates low epistemic uncertainty (tight SHAP concentration), while a flat $\Gamma$ reflects high epistemic uncertainty (broad dispersion). The median SHAP value occurs at $\Gamma(c)=0.5$. We operationalize this framework through uncertainty entropy minimization (Equation: 17). which guides optimal data acquisition (Equation: 18). This entropy reduction disproportionately targets features with high $\text{Var}(\phi_{j})$, thereby reducing aleatoric uncertainty ($\mathbb{E}_{D}[\text{Var}_{f}(\phi_{i}|D)]$). The theory explicitly quantifies epistemic uncertainty through $\Gamma$’s spread, complementing the evidence theory framework (Algorithm: 4).
Dirichlet Process Hypothesis Sampling
Dirichlet process (DP) hypothesis sampling (Algorithm: 1) integrates both aleatoric and epistemic uncertainty through Bayesian nonparametrics. We model the posterior over tree ensembles (Equation:  20). where $G_{0}$ weights trees by out-of-bag reliability. The concentration parameter $\alpha$ controls uncertainty estimation: $\alpha\ll 1$ focuses on high-accuracy trees (low epistemic uncertainty), while $\alpha\gg 1$ enforces uniform weighting (high epistemic uncertainty). SHAP distributions are derived as $\mathbb{F}_{i}(A)$ (Equation: 22). As $K\to\infty$, $\mathbb{F}_{i}$ converges to a Gaussian process (Equation: 24), preserving SHAP linearity. This method captures aleatoric uncertainty through within-sample SHAP variance and epistemic uncertainty through between-sample variance of $\mathbb{E}[\phi_{i}]$, while maintaining entanglement via the DP’s covariance structure.
The three frameworks form an end-to-end workflow that decomposes SHAP variance (Algorithm: 3) (Equation: 1). Evidence theory quantifies epistemic uncertainty and conflict, Liu’s theory models epistemic spread and guides data acquisition, and Dirichlet sampling integrates both through its weighted nonparametric formulation. The interconnection manifests in three key linkages: (1) Conflict detection (evidence theory) flags features for entropy minimization (Liu’s theory); (2) Dirichlet samples generate distributions feeding into $\text{Bel}/\text{Pl}$ and $\Gamma$ calculations; (3) Data acquisition refines $G_{0}$ in the DP base measure. This triad addresses the SHAP uncertainty decomposition as follows: aleatoric uncertainty is measured through BPA dispersion (evidence theory) and within-DP-sample variance; epistemic uncertainty is quantified by $\text{Pl}(A)-\text{Bel}(A)$, $\Gamma$-entropy, and $\alpha$-driven hypothesis sampling; entanglement is preserved via conflict terms $\mathcal{C}_{i}$ and the DP’s covariance structure. The unified methodology (Algorithm: 5) enables granular attribution of uncertainty sources critical for high-stakes interpretability.
4.1 Physical Interpretation
Evidence Theory
  • Belief: Minimum support for SHAP interval
  • Plausibility: Maximum possible support
  • Conflict: $\text{Pl}(A)-\text{Bel}(A)>0$ indicates ambiguous explanations
Uncertainty Distribution
  • $\Gamma(c)=0.5$ at median SHAP value
  • Steep $\Gamma$ $\Rightarrow$ low epistemic uncertainty
  • Flat $\Gamma$ $\Rightarrow$ high epistemic uncertainty
Dirichlet Process
  • $\alpha$ controls “exploration-exploitation” of hypothesis space
  • $w_{k}$ weights represent tree reliability
  • Samples $G$ represent plausible realizations of the model
4.2 Practical Implications
  • Conflicting Explanations: High $\text{Pl}(A)-\text{Bel}(A)$ triggers human verification in critical applications.
  • Data Acquisition: Minimizing $H(\Gamma)$ focuses data collection on high-uncertainty features:
    $\frac{\partial H}{\partial n_{j}}\propto-\text{Var}(\phi_{j})$ (25)
  • Hypothesis Sampling: The Dirichlet concentration parameter $\alpha$ controls uncertainty estimation:
    • $\alpha\approx 1$: Balanced exploration
    • $\alpha<1$: Focus on best-performing trees
    • $\alpha>1$: Uniform uncertainty estimation
Algorithm 1 Dirichlet-Weighted Tree Sampling
Purpose: Generate hypothesis-consistent sub-ensembles
Input: Trained ensemble $\mathcal{M}$, training data $D$, concentration $\alpha$, temperature $\beta$
Output: List of $S$ sub-ensembles
1:function DirichletSample($\mathcal{M}$, $D$, $S$, $\alpha$, $\beta$)
2:  for each tree $T_{k}$ in $\mathcal{M}$ do
3:   Compute OOB accuracy $a_{k}$ using $D$
4:   $w_{k}\leftarrow\exp(\beta\cdot a_{k})/\sum_{j}\exp(\beta\cdot a_{j})$ $\triangleright$ Softmax weighting
5:  end for
6:  for $s=1$ to $S$ do
7:   Draw $\pi\sim\text{Dirichlet}(\alpha\cdot w)$ $\triangleright$ Dirichlet distribution
8:   Sample tree indices $I\sim\text{Categorical}(\pi)$
9:   Construct sub-ensemble $\mathcal{M}_{s}=\{T_{i}\mid i\in I\}$
10:   return $\mathcal{M}_{s}$
11:  end for
12:end function
Algorithm 2 Constrained TreeSHAP Computation
Purpose: Compute SHAP values preserving path dependencies
Input: Sub-ensemble $\mathcal{M}_{s}$, instance $\mathbf{x}$, background data $B$
Output: SHAP vector $\phi$
1:function ConstrainedTreeSHAP($\mathcal{M}_{s}$, $\mathbf{x}$, $B$)
2:  for each tree $T$ in $\mathcal{M}_{s}$ do
3:   $\phi_{T}\leftarrow\text{TreeSHAP}(T,\mathbf{x},B)$ $\triangleright$ Standard TreeSHAP computation
4:  end for
5:  $\phi_{\text{mean}}\leftarrow\text{mean}(\phi_{T}\text{ across trees})$
6:  $\Sigma\leftarrow\text{Covariance}(\phi_{T}\text{ across trees})$ $\triangleright$ Feature covariance matrix
7:  $\phi_{\text{adj}}\leftarrow\phi_{\text{mean}}+0.5\cdot\text{diag}(\Sigma)$ $\triangleright$ Interaction adjustment
8:  return $\phi_{\text{adj}}$
9:end function
Algorithm 3 SHAP Variance Decomposition
Purpose: Quantify uncertainty components
Input: SHAP distributions $\{\Phi_{s}\}_{s=1}^{S}$
Output: Aleatoric, epistemic, entanglement terms
1:function DecomposeVariance($\{\Phi_{s}\}$)
2:  for each feature $i$ do
3:   $\mu_{s}[i]\leftarrow\text{mean}(\Phi_{s}^{i})$ $\triangleright$ Within-sample mean
4:   $\sigma^{2}_{s}[i]\leftarrow\text{variance}(\Phi_{s}^{i})$ $\triangleright$ Within-sample variance
5:   $A[i]\leftarrow\text{mean}(\sigma^{2}_{s}[i])$ $\triangleright$ Aleatoric uncertainty
6:   $E[i]\leftarrow\text{variance}(\mu_{s}[i])$ $\triangleright$ Epistemic uncertainty
7:   $C[i]\leftarrow\text{Covariance}(\mu_{s}[i],\sigma^{2}_{s}[i])$ $\triangleright$ Entanglement term
8:  end for
9:  return $(A,E,C)$
10:end function
Algorithm 4 Uncertainty-Aware SHAP Aggregation
Purpose: Compute final SHAP values with uncertainty metrics
Input: SHAP distributions $\{\Phi_{s}\}$, features $F$
Output: Mean SHAP, uncertainty metrics
1:function AggregateUncertainty($\{\Phi_{s}\}$, $F$)
2:  for each feature $i$ in $F$ do
3:   $\mu[i]\leftarrow\text{mean}(\Phi_{s}^{i})$ $\triangleright$ Mean SHAP value
4:   $\sigma[i]\leftarrow\text{std}(\Phi_{s}^{i})$ $\triangleright$ Standard deviation
5:   $\text{CI}[i]\leftarrow[\text{percentile}(\Phi_{s}^{i},2.5),\text{percentile}(\Phi_{s}^{i},97.5)]$ $\triangleright$ 95% CI
6:   $H[i]\leftarrow\text{Entropy}(\Phi_{s}^{i})$ $\triangleright$ Differential entropy
7:   $\text{SS}[i]\leftarrow P(\text{sign}(\phi)\text{ constant})$ $\triangleright$ Sign stability
8:  end for
9:  return $(\mu,\sigma,\text{CI},H,\text{SS})$
10:end function
Algorithm 5 UbiQTree End-to-End
Purpose: Full uncertainty quantification pipeline
Input: Model $\mathcal{M}$, data $D$, instance $\mathbf{x}$, parameters
Output: SHAP values with uncertainty
1:function E_SHAP($\mathcal{M}$, $D$, $\mathbf{x}$, $S=500$, $\alpha=0.5$, $\beta=5.0$)
2:  $\triangleright$ Step 1: Hypothesis sampling
3:  $\mathcal{M}_{\text{list}}\leftarrow\text{DirichletSample}(\mathcal{M},D,S,\alpha,\beta)$
4:
5:  $\triangleright$ Step 2: SHAP computation
6:  for each $\mathcal{M}_{s}$ in $\mathcal{M}_{\text{list}}$ do
7:   $\Phi_{s}\leftarrow\text{ConstrainedTreeSHAP}(\mathcal{M}_{s},\mathbf{x},D)$
8:   Store $\Phi_{s}$
9:  end for
10:
11:  $\triangleright$ Step 3: Variance decomposition
12:  $(A,E,C)\leftarrow\text{DecomposeVariance}(\{\Phi_{s}\})$
13:
14:  $\triangleright$ Step 4: Uncertainty metrics
15:  $(\mu,\sigma,\text{CI},H,\text{SS})\leftarrow\text{AggregateUncertainty}(\{\Phi_{s}\})$
16:
17:  return mean_shap: $\mu$, std_dev: $\sigma$, ci_95: CI, aleatoric: A, epistemic: E, entanglement: C, entropy: H, sign_stability: $S$
18:end function
5 Results
To study and analyze epistemic uncertainty in these experiments, we implemented our framework. We performed an ensemble-based SHAP analysis for each class in our dataset. Then, we plotted the mean absolute SHAP value for each class in our dataset using the trained model. For this study, we trained an RF classifier across various datasets. In this study, we relied on the absolute SHAP values for each class in the dataset. These values are useful for comparing the relative strength of a feature’s contribution within each class, measuring the uncertainty of a feature’s impact on the class’s output, and identifying features that the model considers decisive for a class, regardless of whether they increase or decrease the class’s logit/probability. Absolute SHAP is particularly well-suited for quantifying uncertainty per class because it allows us to analyze the stability of a feature’s influence on a given class. Furthermore, it allows us to evaluate whether the model consistently demonstrates confidence in the feature’s significance for the class, regardless of its sign. SHAP variance or entropy indicates the robustness of class-specific attribution magnitude across model variants. The $\pm 2\sigma$ (standard deviation) is plotted on the mean absolute SHAP chart. This $\pm 2\sigma$ denotes epistemic uncertainty. Features are then ranked by mean contribution. Relatively wide violin plot indicate considerable variability across different instantiations of sub-ensembles. Higher contributions show that the model consistently relies on this feature. Along with the wide violin plot, they indicate that its impact magnitude is not well understood and that there is a lot of uncertainty about it. The narrow violin plot on the chart represent high-confidence features that contribute to stable predictions. We select the top three features with the highest contributions from each class of the different datasets; however, the user can select as many as required and analyze them. SHAP distribution analysis is performed to evaluate the stability of features that contribute the most across model instances. A high standard deviation suggests that the SHAP values are inconsistent across subtrees or subensembles, indicating uncertainty about each feature’s influence. The SHAP distribution analysis also reveals explanation entropy. High entropy indicates a flat or dispersed distribution of SHAP values, indicating low certainty about the features’ impact. Consistent, peaked SHAP value distributions indicate low entropy. The SHAP distribution visualization helps users understand the directional stability of SHAP values. This measure quantifies the consistency of the sign of the SHAP value, providing insight into whether the SHAP values contribute negatively or positively across sub-models. We group features into three categories based on their directional consistency: high stability (>=90%), moderate stability (>=67%), and low stability (<67%). This allows us to study how the interpretation values vary despite having varying epistemic uncertainty. This provides insight into the validity of the SHAP values in any given model realization. Overall, quantifying epistemic uncertainty and visualizing SHAP magnitude distributions, distributional shapes, entropies, and directional consistencies helps us quantify uncertainty in terms of feature contributions or importance, as calculated using SHAP values. Analyzing the subtrees in ensemble methods helps us simulate posterior samples from the model space. This allows us to focus on uncertainty in the model rather than data noise. We control our methodology using the number of posterior samples, or the $\alpha$ parameter. We select different parameters for each dataset to validate our approach and collect insights. By extracting and explaining predictions using different subsets of trees (or different models in an ensemble), we simulate how the model would behave under slightly different yet still plausible versions of itself. This captures variability due to uncertainty in model specification. SHAP explanations reveal differences via sub-models and show how much confidence can be placed in a specific feature attribution. This makes them sensitive to the model’s structure. The epistemic uncertainty thresholds are chosen based on experiments: 0.05 <= $\sigma$ < 0.1 requires expert-in-the-loop verification; 0.05 <= $\sigma$ < 0.05 is used for automated decisions; and $\sigma$ >= 0.1 suggests model retraining. A mean $\pm 2\sigma$ SHAP chart shows a feature’s average effect and variability. Wide variance across submodels implies high uncertainty. A SHAP kernel density estimate (KDE) along with the confidence interval (CI) plot visualizes the distribution of SHAP values. Multimodality or flatness indicates disagreement between model variants. Finally, quantitative uncertainty metrics, such as SHAP distribution plot which indicates the standard distriution, entropy, and sign stability, offer summary statistics that allow us to directly assess the robustness and stability of a feature’s contribution from the perspective of evidence theory.
Medical Information Mart for Intensive Care III (MIMIC-III) Dataset
(a) No Admission
(b) Very Short Stay
(c) Short Stay
(d) Long Stay
Figure 1: The SHAP summary plot provides a visual representation of the impact of various features on a model’s prediction for the No Admit, Very Short Stay, Short Stay, & Long Stay class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The color of the each bar corresponds to the absolute SHAP value magnitude, and the light blue shaded area indicates the 95% confidence interval($\pm 2\sigma$) of the feature’s impact on the prediction. The plot suggests that NumTransfers, NumNotes, AdmitProcedure are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the No Admit class; NumTransfers, NumNotes, NumDiagnosis are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Very Short Stay class; NumNotes, NumTransfers, AdmitProcedure are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Short Stay class; NumNotes, NumTransfers, NumDiagnositcs are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Long Stay class.
(a) NumTransfers
(b) NumNotes
(c) AdmitProcedure
Figure 2: The distribution of SHAP values for the four most contributing features to the No Admit class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) NumTransfers
(b) NumNotes
(c) NumDiagnosis
Figure 3: The distribution of SHAP values for the four most contributing features to the Very Short Stay class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) NumNotes
(b) NumTransfers
(c) AdmitProcedure
Figure 4: The distribution of SHAP values for the four most contributing features to the Short Stay class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) NumNotes
(b) NumTransfers
(c) NumDiagnosis
Figure 5: The distribution of SHAP values for the four most contributing features to the Long Stay class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
The Medical Information Mart for Intensive Care III (MIMIC-III) [78] is a substantial clinical database comprising detailed health-related data from over 40,000 adult patients admitted to critical care units at a tertiary care hospital between 2001 and 2012. The dataset under consideration is extensive in nature, encompassing a wide range of information pertinent to the subject. This includes demographic data, vital signs, laboratory test results, medications, procedures, clinical notes, imaging reports, and hospital length of stay. The features ’hadm_id’, ’LOSdays’, ’religion’, ’marital_status’, ’ethnicity’ were removed for ethical and privacy reasons and to replicate the real-world scenarios. We utilised this dataset to classify the datapoints into length of stay (LOS). We then, grouped the classes into No Admission, Very Short Stay, Short Stay, and Long Stay categories. No Admission applies to patients who are assessed but not admitted to the hospital. Very Short Stay refers to hospitalizations lasting less than three days, while Short Stay covers stays from three to seven days. Long Stay describes hospitalizations exceeding seven days [79]. The feature descriptions for the features we have discussed are as follows, NumTransfers represents the number of times a patient is transferred within the hospital during their stay. These transfers may occur between different units, such as from the emergency department to the ICU, or between wards. They reflect patient movement within the hospital.NumNotes is the count of clinical notes recorded for a patient during their hospital admission or ICU stay. These notes include documentation from physicians, nurses, and other healthcare providers, capturing clinical observations, treatments, and progress. Admit Procedures is number or list of procedures performed around the time of hospital admission. These procedures could be diagnostic or therapeutic interventions documented in the procedure coding system and reflect initial clinical care. NumDiagnosis is the number of unique diagnoses assigned to the patient during their hospital or ICU stay. Diagnoses are typically coded using International Classification of Diseases (ICD) codes, which provide an overview of the patient’s clinical conditions. gender is the administrative gender identity of the patient, as recorded in hospital records. In MIMIC, this is usually "M" (male) or "F" (female), though other categories may be included depending on the data source. ExpiredHospital is a binary flag indicating whether the patient died during the hospital stay (1 for yes and 0 for no). This captures in-hospital mortality as a key outcome variable.
The RF Classifier was trained with 100 numbers of estimators along with the default parameters following an 80:20 train test stratified split. The reported F1 score on the dataset was 89.0%. The implementation of the proposed framework on the test data set was undertaken to assess the epistemic uncertainty inherent in predictions pertaining to unseen data. The number of samples, denoted by the parameter "number of samples," was set to 100 sub-models or tree subsets. This was done to simulate 100 posterior samples from the model space. The parameter designated as $\alpha$ is employed to modulate the degree to which sub-trees are explored. In the context of model-based search algorithms, the parameter $\alpha$ serves to determine whether the user’s objective is to identify balanced trees, trees with optimal performance, or trees characterized by uniform uncertainty estimation. For this experiment, the value of $\alpha$ is selected to be less than one, with the objective of identifying the trees with the highest performance. For each of the classes, the absolute SHAP values are calculated (see Figure:  1).
[…]
Furthermore, we examined the top three features, their associated epistemic uncertainty, sign stability, and SHAP value distribution, as well as the mean SHAP and 95% SHAP confidence interval. For the No Admit or No Admission category, the epistemic uncertainties for the features NumTransfers, NumNotes, and AdmitProcedures are 0.063, 0.061, and 0.045, respectively (Figure: 2). The sign stability for these features is 99.0%, 72.0%, and 96.0%, respectively. These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. Explanation entropy for the features NumTransfers is considerably low as well, indicated by a uniform distribution of SHAP values, signaling high information certainty about the feature’s impact. The feature AdmitProcedure has high entropy, indicated by skewed SHAP value distributions. NumNotes have sign stability of 72.0%, indicating interpretive inconsistency despite low epistemic uncertainty. For the Very Short Stay category, the epistemic uncertainties for the features NumTransfers, NumNotes, and NumDiagnosis are 0.063, 0.055, and 0.053, respectively. The sign stabilities for these features are 100.0%, 99.0%, and 96.0%, respectively (Figure 3). The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The sign stabilities are very high, indicating interpretive consistency with low epistemic uncertainty. For the Short Stay category, the epistemic uncertainties for the features NumNotes, NumTransfers, and AdmitProcedures are 0.060, 0.038, and 0.048, respectively. The sign stabilities for these features are 84.0%, 85.0%, and 89.0%, respectively (Figure: 4). The standard deviation of epistemic uncertainties is low in the top three features, suggesting consistent SHAP values across the ensemble and reflecting certainty about the features’ influence. Explanation entropy for the features NumTransfer, AdmitProcedures, & NumNotes is considerably high as well, indicated by a non-uniform distribution of SHAP values, signaling high information certainty about the feature’s impact. The sign stabilities are mostly stable, indicating interpretive inconsistency despite low epistemic uncertainty. For the Long Stay category, the epistemic uncertainties for the features NumNotes, NumTransfers, and NumDiagnosis are 0.041, 0.033, and 0.030, respectively. The sign stabilities for these features are 100.0%, 99.0%, and 90.0%, respectively (Figure:  5). The standard deviation of epistemic uncertainties is low in the top three features, suggesting consistent SHAP values across the ensemble and reflecting certainty about the features’ influence. However, a skewed distribution of SHAP values signals low information certainty about the feature’s impact. The features have sign stability, indicating interpretive consistency with low epistemic uncertainty.
Ovarian Cancer Dataset
(a) No Risk
(b) Low Risk
(c) Medium Risk
(d) High Risk
Figure 6: The SHAP summary plot provides a visual representation of the impact of various features on a model’s prediction for the No Risk, Low Risk, Medium Risk, & High Risk class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The color of the each bar corresponds to the absolute SHAP value magnitude, and the light blue shaded area indicates the 95% confidence interval($\pm\sigma$) of the feature’s impact on the prediction. The plot suggests that Smoking, PreviousTreatment, & HormoneTherapy are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the No Risk class; Symptom, PreviousTreatment, & EnhancementPattern are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Low Risk class; Smoking, PreviousTreatment, & HormoneTherapy, are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the Medium Risk class; Smoking, HormoneTherapy, & PreviousTreatment are the most impactful feature, significantly contributing to the high probability of the prediction’s shift toward the High Risk class.
(a) Smoking
(b) PreviousTreatment
(c) HormoneTherapy
Figure 7: The distribution of SHAP values for the four most contributing features to the No Risk class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) Symptom
(b) PreviousTreatment
(c) EnhancementPattern
Figure 8: The distribution of SHAP values for the four most contributing features to the Low Risk class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) Smoking
(b) HormoneTherapy
(c) PreviousTreatment
Figure 9: The distribution of SHAP values for the four most contributing features to the Medium Risk class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) Smoking
(b) PreviousTreatment
(c) HormoneTherapy
Figure 10: The distribution of SHAP values for the four most contributing features to the High Risk class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
The Ovarian Cancer Dataset [80] contains 200,100 patient records collected hourly between January 2019 and December 2024. This highly detailed longitudinal dataset is useful for monitoring ovarian cancer risk and progression. It is designed to support prognostic modeling and progression risk assessment in ovarian cancer patients. This dataset contains 200,100 data points and 34 features. It was used to categorize ovarian cancer in females into risk categories. For each data point, the associated feature, Risk Label, has four classes: No Risk, Low Risk, Medium Risk, and High Risk. No Risk coressponds to no evidence or indication of risk. Low Risk corresponds to minimal probability of an adverse outcome or malignancy. Medium Risk coressponds to moderate chance of risk; requires monitoring or further evaluation. High Risk indicates a high probability of an adverse outcome or malignancy and likely warrants intervention. The description of the features we have discussed in the results are as mentioned further. Symptom coressponds to clinical signs or patient-reported symptoms that are associated with the progression or risk of ovarian cancer and are used to characterize the disease. PreviousTreatment coressponds to information on any medical therapies or interventions the patient underwent before the current assessment, such as chemotherapy, surgery, or radiation therapy. EnhancementPattern is an imaging-derived feature that describes the contrast enhancement patterns observed in diagnostic scans. Smoking is the patient’s smoking history or status is a known risk factor impacting ovarian cancer progression and overall health. HormoneTherapy/HormoneTreatment are the records of hormonal treatments received by the patient, including exogenous hormone administration, which may influence cancer risk or progression. SocioeconomicStatus is a categorical or continuous measure reflecting the patient’s social and economic circumstances, which can affect access to healthcare and outcomes. SNP_Status is a genetic feature indicating the presence or absence of specific single nucleotide polymorphisms (SNPs) related to ovarian cancer susceptibility or progression.
The dataset was highly imbalanced, with the No Risk class having 119,965 data points, the Low Risk class having 40,092 data points, the Medium Risk class having 30,068 data points, and the High Risk class having 9,975 data points. We implemented oversampling to create a class balance using the hybrid resampling technique SMOTETomek [81], which combines oversampling using SMOTE [82] and undersampling using TOMEK link removal. This technique removes noisy and borderline instances after oversampling. The resulting dataset was then used to create an 80:20 training-testing split with stratified sampling. We trained an RF classifier with number of estimators equal to 500 and default parameters on the training set, achieving an F1 score of 58.0% on the test set. Our main focus with this dataset was studying uncertainty and analyzing it when the model cannot learn complex patterns from the dataset. We implemented our framework on the test data to analyze epistemic uncertainty in predictions on unseen data. For the experiment on this dataset, the 500 sub-models or tree subsets parameter for the number of samples was set to simulate 500 posterior samples from the model space. We chose the $\alpha$ parameter to be 1.0 to perform balanced sampling of the sub-ensembles. We calculated the absolute SHAP values for each risk label. For the No Risk class, the top three most important features were identified to be Smoking, PreviousTreatment, & HormoneTherapy(Figure 6). For the Low Risk class, the top three most important features were identified to be Symptom, PreviousTreatment, & EnhancementPattern. For the MediumRisk class the top three most important features were identified to be Smoking, HormoneTherapy, & PreviousTreatment. For the High Risk class the top three most important features were identified to be Smoking, HormoneTherapy, & PreviousTreatment. The results from the absolute mean SHAP values chart (Figure 6) indicate that SHAP feature importance varies considerably among the underlying models. This suggests considerable variability across different model subsets. While the model consistently relies on features such as Symptom, MenstrualHistory, SocioeconomicStatus, HormoneTreatment, etc. , it does so with substantial epistemic ambiguity regarding their precise impact magnitude. In contrast, features such as SNP_Status exhibit low average SHAP values and narrow uncertainty narrow violin plots, reflecting low importance and high confidence in their negligible contribution. This hints at stable but marginal roles in predicting the different classes. This feature has consistent performance across all classes. One major difference observed is in terms of positively or negatively influencing predictions.
We also studied the top three features of this dataset and their associated epistemic uncertainty, sign stability, and SHAP value distribution, as well as the mean SHAP and 95% SHAP confidence interval. For the No Risk category, the epistemic uncertainties for the features No Risk category the epistemic uncertainties for the features Smoking, PreviousTreatment, & HormoneTherapy are 0.050, 0.078, and 0.057, respectively (Figure 7). The sign stabilities for these features are 100.0%, 94.6%, and 100.0%, respectively. These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP distribution is non-uniform and skewed for the Smoking and HormoneTherapy features, signaling high information uncertainty about the feature’s impact. These features have high sign stability, indicating interpretive consistency and low epistemic uncertainty. For the Low Risk category, the epistemic uncertainties for the features Symptom, PreviousTreatment, & EnhancementPattern are 0.016, 0.063, and 0.020, respectively. The sign stabilities for these features are 98.4%, 71.6%, and 86.8%, respectively (Figure: 8). These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP has a non-uniform distribution for the Symptom and EnhancementPattern features, signaling high information uncertainty about the feature’s impact. The features PreviousTreatment and EnhancementPattern have a sign stability of 71.6% and 86.8%, respectively, indicating not-very-high interpretive consistency despite low epistemic uncertainty. The SHAP distribution exhibits flatness, indicating disagreement across the sub-ensembles. For the Medium Risk category, the epistemic uncertainties for the features Smoking, HormoneTherapy, & PreviousTreatment are 0.050, 0.015, and 0.051, respectively. The sign stabilities for these features are 92.0%, 97.2%, and 82.8%, respectively (see Figure: 9). These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP has a non-uniform distribution for the Smoking and HormoneTheraphy features, signaling high information uncertainty about the feature’s impact. The featurePreviousTreatment have a sign stability of 82.8% indicating not-very-high interpretive consistency despite low epistemic uncertainty. The SHAP distribution exhibits flatness, indicating disagreement across the sub-ensembles. For the High Risk category, the epistemic uncertainties for the features Smoking, PreviousTreatment, & HormoneTherapy are 0.026, 0.036, and 0.040, respectively. The sign stabilities for these features are 100%, 75.4%, and 96.6%, respectively (see Figure: 10). These metrics reflect model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. The SHAP distribution is highly skewed for all the features, signaling high information uncertainty about the feature’s impact, along with a relatively dispersed distribution. This reinforces the model’s uncertainty about precise attribution. The features have a very high sign stability of 75.4%, indicating high interpretive consistency with low epistemic uncertainty.
SEER Breast Cancer Dataset
(a) Deceased Patients
(b) Surviving Patients
Figure 11: The SHAP summary plot provides a visual representation of the impact of various features on a model’s prediction for the Deceased & Alive class, incorporating epistemic uncertainty. The violin plots illustrate the distribution of SHAP values for each feature, with individual hypothesis samples represented by gray points. The color of the each bar corresponds to the absolute SHAP value magnitude, and the light blue shaded area indicates the 95% confidence interval($\pm\sigma$) of the feature’s impact on the prediction. The plot suggests that Age, Regional_Node_Examined, & Regional_Node_Positive are the most impactful feature, significantly contributing to the high probability of the prediction’s shift towards the both Deceased & Alive class, although in different directions.
(a) Age
(b) Regional_Node_Examined
(c) Regional_Node_Positive
Figure 12: The distribution of SHAP values for the four most contributing features to the Alive class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
(a) Age
(b) Regional_Node_Examined
(c) Regional_Node_Positive
Figure 13: The distribution of SHAP values for the four most contributing features to the Dead class was examined to further investigate the stability and epistemic uncertainty of the features. The KDE plot shows the distribution of SHAP values collected from different model samples. The red dashed vertical line marks the mean SHAP value, and the shaded region represents the 95% credible interval.
The SEER Breast Cancer Dataset [83] is a comprehensive cancer registry that provides extensive information on breast cancer cases, including patient demographics, tumor characteristics, treatment details, and survival outcomes. The utilization of this method is prevalent in the analysis of treatment outcomes, as it captures comprehensive, population-based longitudinal data. This capability enables researchers and clinicians to assess the efficacy of diverse interventions across a range of patient populations and clinical settings. For instance, advanced predictive models developed using SEER data apply machine learning to predict individual patient survival and treatment response, enabling personalized treatment decisions and optimizing treatment strategies to extend survival while minimizing adverse effects. This capability is of particular significance in complex cases, such as metastatic breast cancer, where the guidance provided by clinical trials is limited and SEER-based models facilitate decision-making by simulating the outcomes of various treatment options that are tailored to the unique characteristics of each patient and tumor [84, 85]. The dataset contains 4,024 data points, 16 features, and associated survival outcomes, specifically the Alive and Dead status, for female patients diagnosed with breast cancer.
The Alive class indicates the patients who were alive at the final follow-up or censoring time. Dead class indicates patients who died from breast cancer or any other cause during the observation period. The features under discussion are described further text. Age: The age of the patient at the time of the initial breast cancer diagnosis. This influences prognosis and treatment choice. Regional_Node_Positive: The number of regional lymph nodes confirmed positive for cancer, indicating spread. Grade_Poorly differentiated (Grade III): Tumor cells are highly abnormal and grow/spread aggressively. Grade_Moderately differentiated (Grade II): Tumor cells are highly abnormal and grow/spread aggressively. Regional_Node_Examined indicating that the total number of regional lymph nodes examined for cancer involvement. Progesterone_Status_Negative indicating that the tumor cells lack progesterone receptors, which may affect response to hormone therapy. T_Stage_T4 corresponding that the tumor has invaded the chest wall and/or skin (advanced tumor size/invasion stage). N_Stage_N1 denoting that the cancer has spread to one to three axillary lymph nodes. N_Stage_N2 indicating that the cancer has spread to four to nine axillary lymph nodes, or the nodes are fixed or matted.N_Stage_N3 indicating that the cancer has spread to 10 or more axillary nodes, as well as to the infraclavicular or internal mammary nodes. 6th_Stage_IIA is the AJCC 6th edition Stage IIA indicating moderately advanced local disease.6th_Stage_IIB is the AJCC 6th edition Stage IIB indicating larger tumor size and/or more node involvement.6th_Stage_IIIA is the AJCC 6th edition Stage IIIA indicating advanced local spread to several regional nodes.6th_Stage_IIIB is the AJCC 6th edition Stage IIIB indicating tumor involves chest wall or skin and may involve nodes.6th_Stage_IIIC is the AJCC 6th edition Stage IIIC indicating extensive lymph node involvement near collarbone or breastbone. The dataset was then employed for the purpose of survival classification to study the treatment outcome. The 80:20 stratified train-test split was implemented. Subsequently, an RF (random forest) classifier with a number of estimators equal to 300 along with default parameters, was trained. The model was evaluated on the test set, and the resultant F1 score was 79.8%. The implementation of the proposed framework on the test data set was undertaken to assess the epistemic uncertainty inherent in predictions pertaining to test data. The value of the $\alpha$ parameter was set to 0.5 to perform the balanced exploration of the trees.
The absolute mean SHAP values chart (Figure: 11) shows that the SHAP feature importance varies considerably among the top contributing features in the underlying models. This suggests considerable variability across different model sub-ensembles. While the model consistently relies on features such as Age, Regional_Node_Positive, Grade_Poorly differentiated: Grade III, Grade_Moderately differentiated: Grade II, Regional_Node_Examined, Progesterone_Status_Negative, the model exhibits substantial epistemic ambiguity regarding their precise impact magnitude. In contrast, features such as T_Stage_T4, N_Stage_N1, N_Stage_N2, N_Stage_N3, 6th_Stage_IIA, 6th_Stage_IIB, 6th_Stage_IIIA, 6th_Stage_IIIB, 6th_Stage_IIIC exhibit low average SHAP values and narrow uncertainty violin plots. This reflects both low importance and high confidence in their negligible contribution, hinting at stable but marginal roles in predicting the different classes. One major difference observed is in terms of positively or negatively influencing predictions. For each survival status, we calculate the absolute SHAP values. For the Dead or Deceased class, the top three most important features were identified as Age, Regional_Node_Examined, & Regional_Node_Positive. For the Surviving or Alive class, the top three most important features are Age, Regional_Node_Examined, & Regional_Node_Positive (Figure: 11). We also examined the top three features of this dataset and their associated epistemic uncertainty, sign stability, and SHAP value distribution with mean SHAP and 95% SHAP confidence intervals. For the Surviving or Alive category, the epistemic uncertainties for the features Age, Regional_Node_Examined, & Regional_Node_Positive are 0.042, 0.044, and 0.037, respectively. The sign stabilities for these features are 98.3%, 94.7%, and 92.0%, respectively (Figure: 12). For the Deceased or Dead category, the epistemic uncertainties for the features Age, Regional_Node_Examined, &
Regional_Node_Positive are 0.042, 0.044, and 0.037, respectively (Figure  13). The sign stabilities for these features are 98.3%, 94.7%, and 92.0%, respectively. For both classes, the metric reflects model variance in attribution for each feature. The low standard deviation of epistemic uncertainties in the top three features suggests consistent SHAP values across the ensemble and reflects certainty about the features’ influence. SHAP values are non-uniform for Age, Regional_Node_Examined, & Regional_Node_Positive features, signaling high uncertainty about the features’ impact, along with a relatively dispersed distribution. This reinforces the model’s uncertainty about precise attribution. The sign stabilities indicate high consistency with low epistemic uncertainty.
6 Conclusion
This research study decomposes SHAP into two categories of uncertainty quantification: aleatoric and epistemic. Breaking down uncertainty allowed us to determine whether it arises from uneliminatable noise in the data or from a lack of knowledge about the true data/model distribution. We also introduced an entanglement term that captures the interaction or covariance between data and model uncertainties. Our approach provides a deeper understanding of SHAP uncertainty in terms of intervals and enables investigation of its origin. In domains such as healthcare, where the consequences can be significant, our method is useful. The proposed approach captures uncertainties that simple intervals cannot while aligning with modern uncertainty quantification practices. DST quantifies both certainty ($Bel$) and possibility ($Pl$) for SHAP attributions, allowing for the explicit modeling of ignorance and epistemic uncertainty. This is particularly useful in ensemble bagging models with conflicting feature attributions. DST offers a non-probabilistic method of expressing confidence in SHAP values, which is useful when data is scarce or evidence is subjective. The framework extends beyond classical probability, offering tools for interpreting and managing uncertainty in model explanations. Identifying constrained SHAP intervals is often challenging in practice for large or complex models [86]. The proposed framework aims to facilitate the estimation of simple uncertainty intervals by leveraging the structural properties of belief functions, uncertainty theorems, and Dirichlet processes. These processes can be efficiently sampled or approximated, thereby enhancing the framework’s efficacy and accessibility. The framework provides control parameters, such as $\alpha$, which enable users to control the exploration-exploitation process of the hypothesis space. Users can choose balanced exploration, exploration of the best-performing trees, or uniform uncertainty estimation. Furthermore, uncertainty theory quantifies confidence in attribution magnitude, with entropy minimization guiding optimal data acquisition. This could facilitate explanations of SHAP values that account for uncertainty, even in high-dimensional or real-world settings. Additionally, our framework enhances uncertainty reporting. In a healthcare context, for example, SHAP values with 95% confidence intervals replace point estimates. The confidence-triggered verification is essential for determining which features require a domain expert’s review. Features with $phi_{i}<0.8$ require a review from a domain expert. This work also touches on the boundaries of AI regulations which emphasizes on the fact that the AI frameworks must provide auditable uncertainty metrics [87]. Together, variance, entropy, and sign stability provide a complete picture of uncertainty. Each chart provides insight into the stability of a SHAP attribution. The user can examine standard deviation, entropy, and sign stability. A mean $\pm 2\sigma$ SHAP indicates high epistemic uncertainty if SHAP varies widely across sub-models. The SHAP confidence interval represents the distribution of SHAP values; multimodality or flatness indicates disagreement across model variants. Uncertainty metrics quantify the reliability and stability of a feature’s attribution.
Acknowledgments
One of the authors of this work has been financially supported by the German Federal Ministry of Health (BMG) under grant No.: ZMI5- 2523GHP027 (project “Strengthening National Immunization Technical Advisory Groups and their Evidence-based Decision-making in the WHO European Region and Globally” SENSE) part of the Global Health Protection Programme, GHPP.